# Adaptive In-conversation Team Building for Language Model Agents

### v2

@@ 3.3 Main Results
| Two-Agents | 3.10 | 2.82 | 5.32 | 1.34 | 2.33 | 5.2 | |
| Meta-prompting | 2.92 | 9.88 | 8.64 | 4.18 | 4.96 | 5.8 | |
| AutoAgent | 4.59 | 18.32 | 33.58 | 12.48 | 12.28 | 7 | |
| DyLAN | 3.01 | 8.76 | - | 7.10 | 8.07 | - | |
| AgentVerse | 7.63 | 13.59 | - | 26.34 | 23.56 | - | |
@@
| Nested Chat Agent Backbone | Mathematics | Programming | Data Analysis | (Sci) Chemistry | (Sci) Physics |
| Blackbox Models | |||||
[…]
| w/ claude-3-sonnet | 35.29 | 80.00 | 60.00 | 15.79 | 26.32 |
[…]
| Open-weight Models | |||||
[…]
| w/ Mixtral-8x22B-instruct-v0.1 | 29.41 | 76.00 | 55.00 | 47.37 | 21.05 |
Table 2 and 3 report the comparison results between Captain Agent and eight different baselines on six real-world scenarios. Baseline results on world information retrieval are extracted directly from the GAIA leaderboard.
Findings 2: Adaptive team-building boosts performance with no task preference. It is obvious that Captain Agent achieves outstanding results over all scenarios, indicating that Captain Agent is free from task preference. Incorporating different agents into the team at a proper time gives Captain Agent the ability to solve difficult tasks like science and world-information retrieval problems step-by-step. On the other hand, Meta-prompting fails in science scenarios due to the inability to decompose science problems into the fine-grain subtasks that one agent can solve. Captain Agent with the agent-team building paradigm neither requires a task that can be decomposed into a subtask that can only be solved by an agent nor requires all agents to be involved in the conversation. We further discuss the static and adaptive teams in Section 3.4.1.
3.4 Analysis and Ablation Studies
3.4.1 Static vs. adaptive team-building
To further explore the power of adaptive team-building, we compare adaptive team-building with static team-building. Specifically, we perform a task-specific team-building paradigm by building a team of agents in the same way as Captain Agent at the beginning of each task and letting them solve each problem. We summarized the results in Table 4, showing that the adaptive team-building paradigm outperforms the static team-building paradigm comprehensively.
3.4.2 Ablation on tool library and agent library
In this part, we conduct an ablation study on the utility of tool and agent libraries. We remove the tool library, the agent library, and both libraries in turn and evaluate the performance on world-information retrieval tasks, i.e., the GAIA dataset. As shown in Table 5, removing the agent library and tool library can both significantly impair the system’s performance. While both the tool and agent libraries can enhance performance independently, optimal results are achieved only when both libraries are employed concurrently. Handling level 1 tasks requires a moderate amount of web browsing and reasoning steps, which can be achieved by several single-turn tool calls or experts writing and executing code iteratively. Introducing both an agent library and tool library makes the system more stable and robust to unknown errors during web interaction, therefore improving the performance. Notably, without an agent library, Captain Agent performs much worse on Level 2 tasks. This is because these tasks are more sophisticated and mostly involve a significant number of web navigation and reasoning steps. Web browsing involves complex and dynamic interactions that are poorly suited to static tool libraries. The tasks require agents to coordinate multiple tools to solve them, which is a process prone to error in web scenarios filled with uncertainty.
3.4.3 Ablation on LLM Backbone
In this section, we explore the influence of the choice of backbone LLM on the performance of Captain Agent. We conduct two experiment settings: weak LLM for Captain Agent and team members, strong backbone for Captain Agent, and weak LLM for nested chat members.
We first equip Captain Agent and its nested experts with four different backbones, namely gpt-4-0125-preview, gpt-4o-mini, LLaMA-3-70B-Instruct, and LLaMA-3-8B- Instruct, and compare it with all the baselines equipped with gpt-4-0125-preview. As shown in Table 6, Captain Agent with gpt-4o-mini outperforms all other baselines.
We then fix the backbone of Captain Agent to gpt-4-0125-preview and employ different backbone LLM for the experts in nested chat, including gpt-3.5-turbo, claude-3-sonnet, gemini-1.5-pro, and open-weight models like LLaMA-3-70B and Mixtral-8x22B. We record the results in Table 7. Chat members with gemini-1.5-pro performs second best in most scenarios. When comparing the results of the two settings, we observe that by utilizing a stronger LLM backbone in Captain Agent to guide the nested conversation, the system’s performance is significantly enhanced.
3.4.4 Cost Analysis
The high token cost associated with LLMs has always been a significant barrier in the practical deployment of agents, rendering them economically unfeasible. We calculate the whole cost of Captain Agent workflow, including generating Captain Agent output, performing agent and tool selection, expert generation, and nested chat conversation. The cost is reported in Table 6. By leveraging smaller, more cost-efficient gpt-4o-mini, our approach significantly reduces costs while maintaining strong performance, achieving an average cost as low as $0.33 per task.
4 Related Work
Large language models (LLMs) represent a significant advancement in artificial intelligence, showcasing remarkable capabilities in various aspects, including reasoning (Wei et al., 2022b; Yao et al., 2024; Morishita et al., 2023; Zhang et al., 2023b; Li et al., 2023a; Ho et al., 2022), planning (BabyAGI, 2023; Song et al., 2023; Valmeekam et al., 2023; Liu et al., 2023b), and adaptability to novel real-world observations (Shi et al., 2024; Hong et al., 2023; Yang et al., 2023a; Dan et al., 2023; Zhou et al., 2023a; Bharadhwaj et al., 2023). Leveraging the inherent versatility of LLMs as generalized models adaptable to diverse scenarios, numerous efforts have been dedicated to the development of intelligent agents (Wu et al., 2023; Xi et al., 2023; Zhang et al., 2024b; Sumers et al., 2023; Zhou et al., 2023b) where LLMs serve as foundational components. For instance, one typical algorithm, React (Yao et al., 2022), employs one single LLM to iteratively generate both reasoning trajectories and task-specific actions. This interleaved process enables the agent to engage in dynamic reasoning. In addition, LLM agents can also harness external tools (Qin et al., 2023a; b; Schick et al., 2024; Cai et al., 2023; Yuan et al., 2023a; Paranjape et al., 2023; Zhang et al., 2024b; Huang et al., 2023; Ma et al., 2024), leveraging both their internal capabilities and external resources, collaborating effectively to solve more intricate problems.
The success of a single-agent system motivates the development of multiple-agent systems (Wang et al., 2023a; Xi et al., 2023; Chen et al., 2023; Wu et al., 2023; Suzgun & Kalai, 2024a; Hong et al., 2023; Zhang et al., 2024b; 2023a; Valmeekam et al., 2023; Wang et al., 2024; Saha et al., 2023; Liang et al., 2023; Du et al., 2023). Methods focusing on static build require a protocol for agents to communicate with each other in a group chat and a builder that can receive the user’s instruction and output an agent list (Wu et al., 2023; Chen et al., 2023; Hong et al., 2023). The builder can be a human (Wu et al., 2023; Hong et al., 2023) or a LLM agent (Chen et al., 2023). There are other works breaking down complex tasks into smaller components, each of which is then handled by a single specialized agent with detailed natural-language instructions (Suzgun & Kalai, 2024b; Zhuge et al., 2023). This task decomposition reduces the prediction burden on each agent by avoiding irrelevant context. For instance, meta-prompting (Suzgun & Kalai, 2024b) involves a meta-model decomposing tasks and assigning subtasks to different LLMs for completion and aggregation.
5 Conclusion and Discussion
We introduce a new paradigm for multi-agent team-building, adaptive build. This new paradigm helps ensure diversity, prevent limited knowledge extraction and reduce stereotypical outputs. The new paradigm executed by our proposed agent, Captain Agent, manages agent teams for problem-solving steps using adaptive multi-agent team building and nested group conversation and reflection. Experimental results across six real-world scenarios demonstrate Captain Agent’s efficacy in various tasks without prompt engineering, achieving superior results compared to existing methods. Ablation studies confirm that each component contributes equally to overall performance, underscoring the robustness of our approach.
Appendix A Limitations
The first limitation of our work is cost. A conversation involving the GPT-4 model costs more than a single-agent system. Although we have reduced the cost by decreasing the participant nested group chat agents, it still has a large conversation and profile as context input. The trade-off between performance and cost will become one of the possible future works for further exploration, like window context, conversation pruning, or conversation compression. Another limitation of our work is the lack of thinking about model diversity. In Table 7, we have demonstrated that the model has task preference, which will influence the nested chat quality. However, before we go deep into the discussion of model preference, we should also notice that the current evaluation of LLM is not perfect. Data leaking is widespread in the pertaining process and will cause the misalignment between the test and real-world performance (Zhang et al., 2024a; Xu et al., 2024). Therefore, a comprehensive yet fair evaluation is important for us to further discuss the ability of model diversity.
Appendix B Social Impact
Appendix C Difference Between Other Team-building Frameworks
@@ Appendix E Task Instructions
We design instructions manually for each scenario and ensure all baselines and Captain Agent receive the same instructions for a fair comparison333Except for the world information retrieval scenario (GAIA), in which we use the results directly from the leaderboard.. All instructions include the basic information of the scenario and may suggest some possible Python libraries, including pandas, numpy, scipy, and sympy.
@@ E.5 Instruction for World-information Retreival
Appendix F Case Studies
Figure 4 illustrates the free-form tool-using ability in the nested conversation when solving a problem in GAIA. Four agents involved in the conversation: DigitalMdeia_Expert, Ornithology_Expert, VideoContentAnalysis_Expert, and UserProxy, in which DigitalMdeia_Expert use perform_web_search tools to request the result of "BBC Earth YouTube Top 5 Silliest Animal Moments" from internet, and VideoContentAnalysis_Expert use get_youtube_subtitle tool to seek for the subtitle from a specific video. After their collaboration, they successfully obtained a correct answer, "Rockhopper penguin."
Appendix G Agent Library
Our agent library recorded 541 agents, including 540 generated agents and one hand-crafted ConversableAgent archived in AutoGen (WebSurferAgent). Here is an example of the agent recorded in the agent library:
We also summarized the agent-selected rate for each scenario in Figure 5. It is obvious that selected agents are highly related to the current scenarios. The verification expert has a high selection rate because we prompt Captain Agent in the system message to create a verification role to maintain the conversation. We also notice that in some specific scenarios (mathematics, data analysis, and programming), some agents with a general name and description will have a high selection rate (e.g., PythonMath_Expert, DataAnalysis_Expert, CodeReview_Expert, etc.). However, in the Science scenarios, there are no highly preferred agents with general descriptions, and the selection distribution become flatten.
Appendix H Tool Library

### v3

@@ 3.4 Analysis and Ablation Studies
3.4.2 Tool library and agent library
To demonstrate the effectiveness of agent and tool library, we conduct ablation studies over three settings: (1) remove both agent and tool library, (2) remove agent library while retaining the tool library, (3) remove tool library while retaining agent library, and (4) keep both agent and tool libraries. We selected GAIA as our target dataset because it represents the strong requirement for high-quality tools and the diversity of problem-solving abilities in different domains, which perfectly match the purpose of this ablation study. As shown in Table 4, removing the agent library and tool library can both significantly impair the system’s performance. While both the tool and agent libraries can enhance performance independently, optimal results are achieved only when both libraries are employed concurrently. Handling level 1 tasks requires a moderate amount of web browsing and reasoning steps, which can be achieved by several single-turn tool calls or experts writing and executing code iteratively. Notably, without an agent library, Captain Agent performs much worse on Level 2 tasks. This is because these tasks are more sophisticated and mostly involve a significant number of web navigation and reasoning steps. Web browsing involves complex and dynamic interactions that are poorly suited to static tool libraries. The tasks require agents to coordinate multiple tools to solve them, which is a process prone to error in web scenarios filled with uncertainty.
3.4.3 Ablation on Reflector
Reflection is one of the most important components in Captain Agent as its feedback represents the essence of the whole nested conversation and will, therefore, largely impact the decision-making process of Captain Agent. In Figure 3, we statistic the performance of Captain Agent with and without reflection and the success rate of the reflector. We define the success rate as: if the LLM-based reflector can detect the dispute in nested conversation and highlight it for Captain Agent to avoid potential next-step error, we will mark it as ”success.” We calculate the success rate of the reflector across all problems under the corresponding scenario. We observe that in (sci) physics problems, the reflector represented a vital role with a high success rate and helped improve the performance of Captain Agent, but the performance of Captain Agent under physics remains lowest across all other scenarios. This is because the reflector cannot refine the problem-solving process but can detect the conflict inside a nested conversation and infer the possible error that may be incurred by the conflict. Therefore, the reflection mechanism helps prevent error propagation in the multi-round problem-solving process designed for Captain Agent.
3.4.4 Ablation on LLM Backbone and Cost Analysis
The high token cost associated with LLMs has always been a significant barrier to the practical deployment of agents, rendering them economically unfeasible. However, with the development of model distillation, the small model can also achieve competitive results with an agent framework on complex tasks. To evaluate Captain Agent’s ability to generalize to different LLM backbones for cost reduction, we equip Captain Agent and its nested experts with four different backbones, including gpt-4-0125-preview, gpt-4o-mini, LLaMA-3-70B-Instruct, and LLaMA-3-8B-Instruct, and compare it with all the baselines equipped with gpt-4-0125-preview. Our cost statistic covers the whole Captain Agent workflow, including Captain Agent output, performing agent and tool selection, agent generation, and nested group conversation. We summarize the results based on the ranking of each setting in Figure 4. We observe that (1) Captain Agent with the same backbone as baselines can achieve the best performance but also the highest cost; (2) Captain Agent with gpt-4o-mini outperforms all other baselines but with a significantly lower cost, revealing the possibility of using a small language model as the backbone of an agent framework without seeing cost as a major issue.
3.4.5 Agent Selected in Each Scenario
Our agent library records 541 agents in total after we complete all experiments, and we summarized the top-10 selected agent’s frequency for each scenario in Figure 5. It is obvious that selected agents are highly related to the corresponding scenarios because we prompt Captain Agent in the system message to create a verification role to keep the conversation going smoothly. We also notice that in some scenarios (mathematics, data analysis, and programming), agents with a non-specific name and description will have a high selection rate (e.g., PythonMath_Expert, DataAnalysis_Expert, CodeReview_Expert, etc.), as the embedding similarity between their description and user task is generally high. But we can still observe some specific agents being selected, like AppliedMath_Expert, StringManipulation_Expert, and NumPySciPy_Expert, etc. For the information retrieval scenario, WebResearch_Expert has an extremely high selection rate. This agent is one of our hand-crafted agents who can search the internet and summarize a page in Markdown format, and Agent Filter can select this agent with Captain Agent instruction properly.
4 Related Work
Large language models (LLMs) represent a significant advancement in artificial intelligence, showcasing remarkable capabilities in various aspects, including reasoning (Wei et al., 2022b; Yao et al., 2024; Morishita et al., 2023; Zhang et al., 2023b; Li et al., 2023a; Ho et al., 2022), planning (BabyAGI, 2023; Song et al., 2023; Valmeekam et al., 2023; Liu et al., 2023b), and adaptability to novel real-world observations (Shi et al., 2024; Hong et al., 2023; Yang et al., 2023a; Dan et al., 2023; Zhou et al., 2023a; Bharadhwaj et al., 2023). Leveraging the inherent versatility of LLMs as generalized models adaptable to diverse scenarios, numerous efforts have been dedicated to the development of intelligent agents (Wu et al., 2023; Xi et al., 2023; Zhang et al., 2024b; Sumers et al., 2023; Zhou et al., 2023b) where LLMs serve as foundational components. For instance, one typical algorithm, React (Yao et al., 2022), employs one single LLM to iteratively generate both reasoning trajectories and task-specific actions. This interleaved process enables the agent to engage in dynamic reasoning. In addition, LLM agents can also harness external tools (Qin et al., 2023a, b; Schick et al., 2024; Cai et al., 2023; Yuan et al., 2023a; Paranjape et al., 2023; Zhang et al., 2024b; Huang et al., 2023; Ma et al., 2024), leveraging both their internal capabilities and external resources, collaborating effectively to solve more intricate problems.
The success of a single-agent system motivates the development of multiple-agent systems (Wang et al., 2023a; Xi et al., 2023; Chen et al., 2023; Wu et al., 2023; Suzgun & Kalai, 2024a; Hong et al., 2023; Zhang et al., 2024b, 2023a; Valmeekam et al., 2023; Wang et al., 2024; Saha et al., 2023; Liang et al., 2023; Du et al., 2023). Methods focusing on static build require a protocol for agents to communicate with each other in a group chat and a builder that can receive the user’s instruction and output an agent list (Wu et al., 2023; Chen et al., 2023; Hong et al., 2023). The builder can be a human (Wu et al., 2023; Hong et al., 2023) or a LLM agent (Chen et al., 2023). There are other works breaking down complex tasks into smaller components, each of which is then handled by a single specialized agent with detailed natural-language instructions (Suzgun & Kalai, 2024b; Zhuge et al., 2023). This task decomposition reduces the prediction burden on each agent by avoiding irrelevant context. For instance, meta-prompting (Suzgun & Kalai, 2024b) involves a meta-model decomposing tasks and assigning subtasks to different LLMs for completion and aggregation. However, these methods lack a backtracking mechanism, making the problem-solving prototype fixed once agents are generated.
5 Conclusion and Discussion
We introduce a new paradigm for multi-agent team-building, adaptive build. This new paradigm helps ensure diversity and prevent limited knowledge extraction. The new paradigm executed by our proposed agent, Captain Agent, manages agent teams for problem-solving steps using adaptive multi-agent team building and nested group conversation and reflection. Experimental results across six real-world scenarios demonstrate Captain Agent’s efficacy in various tasks without prompt engineering, achieving superior results compared to existing methods. Ablation studies confirm that each component contributes equally to overall performance, underscoring the robustness of our approach.
Impact Statement
Appendix A Limitations
The first limitation of our work is cost. A conversation involving the GPT-4 model costs more than a single-agent system. Although we have reduced the cost by decreasing the participant nested group chat agents, it still has a large conversation and profile as context input. The trade-off between performance and cost will become one of the possible future works for further exploration, like window context, conversation pruning, or conversation compression. Another limitation of our work is the lack of thinking about model selection. We should also notice that the current evaluation of LLM is not perfect. Data leaking is widespread in the pertaining process and will cause the misalignment between the test and real-world performance (Zhang et al., 2024a; Xu et al., 2024). Therefore, a comprehensive yet fair evaluation is important for us to further discuss the ability of model diversity. In this work, all agents are communicated under AutoGen with a highly customized communication framework. This limited us to generalize Captain Agent to other existing hand-crafted agents (i.e., absorb those fantastic agents into our agent library). Therefore, a universal agent communication protocol is important.
Appendix B Difference Between Existing Team-building Frameworks
@@
| Two-Agents (gpt-4-0125-preview) | 3.10 | 2.82 | 5.32 | 1.34 | 2.33 | 5.2 | |
| Meta-prompting (gpt-4-0125-preview) | 2.92 | 9.88 | 8.64 | 4.18 | 4.96 | 5.8 | |
| AutoAgent (gpt-4-0125-preview) | 4.59 | 18.32 | 33.58 | 12.48 | 12.28 | 7 | |
| DyLAN (gpt-4-0125-preview) | 3.01 | 8.76 | - | 7.10 | 8.07 | - | |
| AgentVerse (gpt-4-0125-preview) | 7.63 | 13.59 | - | 26.34 | 23.56 | - | |
@@
Appendix C Case Study Comparing Captain Agent and Others
We demonstrate the benefits and functionalities of Captain Agent in two distinct scenarios. The first scenario involves the automatic tailoring of specialized agentic systems for solving complex user queries. The second scenario highlights its ability to replace manually designed agent systems without human intervention, achieving similar performance to the handcrafted three-agent system OptiGuide (Li et al., 2023b) in a specific supply chain optimization scenario.
C.1 Case Study 1: Arbitrary User Question Answering
As illustrated in Figure 6, a user seeks specific information from an article by Emily Midkiff, published in June 2014. In response, Captain Agent assembles a team of specialized agents: WebBrowsing_Expert, Myth_Expert, and Verification_Expert. Each agent is designed to address a specific domain, and together they effectively collaborate to solve the problem, under the coordination of Captain Agent. In contrast, a simple two-agent system from AutoGen fails to resolve the issue due to limitations in the Assistant’s capabilities. This highlights Captain Agent’s ability to build an optimal team of agents and ensure their collaboration, whereas basic two-agent systems may lack the necessary functionalities to address more complex queries.
C.2 Case Study 2: Domain Specific Question Answering
Figure 7 illustrates a comparison between Captain Agent and OptiGuide (Li et al., 2023b), a handcrafted three-agent system designed for supply chain optimization. When provided with the necessary background information, Captain Agent is able to generate a similar group structure, successfully solving the task. In contrast, an examination of the implementation code reveals that OptiGuide includes a fixed structure requiring over 100 lines of code, meticulously crafted by humans. In comparison, CaptainAgent accomplishes the same task with only a few lines of code, significantly reducing the complexity for developers and easing domain-specific development.
@@ Appendix E Task Instructions
We design instructions manually for each scenario and ensure all baselines and Captain Agent receive the same instructions for a fair comparison333Except for the information retrieval scenario (GAIA), in which we use the results directly from the leaderboard.. All instructions include the basic information of the scenario and may suggest some possible Python libraries, including pandas, numpy, scipy, and sympy.
@@ E.5 Instruction for World-information Retreival
Appendix F Agent Library
We provide an example of the agent recorded in the agent library below:
Appendix G Tool Library